OpenStudy (anonymous):

Please Help Me Out!! The length of a rectangle is twice its width. If the area of the rectangle is 128 yd^2, find its perimeter.

OpenStudy (carson889):

Knowns: Area = length * width Perimeter = 2*length + 2*width Length = 2* width Area = 128 yd^2 Unknowns: Perimeter. --------------------- Perimeter = 2*l + 2*w Since length = 2*width Perimeter = 4*w + 2*w = 6*w Now, Area = length * width = 2*width * width = 2*width^2 Rearrange to solve for width: A = 2*w^2 A/2 = w^2 sqrt(A/2) = w We know Area (A) is 128, so after subbing in that value we get: w = sqrt(128/2) = 8 Now insert the formula for width into our refined formula for perimeter P = 6*w P = 6*8 = 48 yd Therefore, the perimeter is 48 yds.